OpenStudy (anonymous):

. In 2010, the population of a city was 65,903. It was expected to grow at a rate of 0.7% each year. What is best prediction for the city’s population in 2027? 66,364 74,200 82,119 112,035

OpenStudy (anonymous):

@iGreen Please Help

OpenStudy (anonymous):

oh are you done

OpenStudy (igreen):

I'm thinking it grows at the same rate of 0.7%? We can write the exponential function: \(f(x) = 65903(1.007)^x\) where x is years. Plug in 10 and solve, what do you get?

OpenStudy (igreen):

Actually plug in 17..my bad..'cause 2027 - 2010 = 17

OpenStudy (igreen):

\(f(x) = 65903(1.007)^{17}\) First simplify \(1.007^{17}\) then multiply that by 65,903.

OpenStudy (anonymous):

f(x)= 66364.321

OpenStudy (anonymous):

is that it

OpenStudy (igreen):

No..first do 1.007^17, what do you get?

OpenStudy (igreen):

@Xgloves11